If I have an isosceles right triangle with the longest side having 48", what is the length of the 2 remaining sides? Please show the formula.....thank you very much.
Answer by fractalier(6550)

Well the Pythagorean Theorem says
a^2 + b^2 = c^2
but since the triangle is isosceles, a = b, so we have
a^2 + a^2 = 48^2
2a^2 = 2304
a^2 = 1152
a = 24*sqrt(2)